find an equation of the line satisfying the conditions given. express your answer in the standard form.
parallel to 5x - y =4 and passing through (-2, 0)

Answer by mananth(16946)

5 x + -1 y = 4
Find the slope of this line
make y the subject
-1 y = -5 x + 4
Divide by -1
y = 5 x -4
Compare this equation with y=mx+b
slope m = 5

m= 5 ,point ( -2 , 0 )
Find b by plugging the values of m & the point in
y=mx+b
0 = -10.00 + b
b= 10
m= 5
Plug value of the slope and b in y = mx +b
The required equation is y = 5 x + 10
y = 5 x + 10
-5x+y=10
m.ananth@hotmail.ca